# RAPTOR: A Foundation Policy for Quadrotor Control

### v1

@@ 2.4.2 Outdoor Tests
2.4.3 Disturbances
Figure 6: RAPTOR Policy in Different Situations. (A): The foundation policy is activated in mid-air, starting with a linear velocity of $4.5$ m/s. (B) Crazyflie Brushless tracking an agile trajectory. (C) Long exposure photo of three consecutive loops of the Crazyflie Brushless trajectory. (D) The foundation policy recovers from being poked and tilting $90^{\circ}$. (E) A tool is rested onto the quadrotor flown by the foundation policy. (F) A quadrotor with four different propellers (2- and 3-blade) is tracking a trajectory using the foundation policy.
@@ 2.4.3 Disturbances
Additionally, we also test swapping out 1, 2, and 3 propellers with random three-blade ones (the original ones are two-blade) and find that, in any of these configurations, the foundation policy still stably controls the vehicle and is even able to track trajectories (cf. Figure 6F and Movie S6). Note that the distribution of quadrotors described in Section 4.1 covers only quadrotors with identical thrust curves for all four motors/propellers. Hence, neither the teacher policies nor the foundation policy has ever experienced differing thrust curves on the same quadrotor. Yet, we find that the foundation policy can generalize zero-shot outside of this distribution, and control the quadrotor well.
In the absence of disturbances, we find that the foundation policy yields a remarkably repeatable performance, as can be seen from the trajectories in Figure 5, which show $5$ consecutive full loops each. This also answers the posed research question about the trajectory length generalization of the foundation policy. During Meta-Imitation Learning, we train the policy with sequences of $500$ steps, corresponding to $5$ s of flight time. During inference, we test the foundation policy with $5$ consecutive iterations of, e.g., a $10$ s, amounting to a trajectory length of $50$ s. This shows a $10\times$ context window size extrapolation.
@@ 2.4.4 Simulation
To further test the out-of-distribution generalization beyond the aforementioned real-world tests including flexible frame, mixed propellers, context window extrapolation, etc., we test simulation-to-simulation transfer to the Flightmare simulator. This transfer is interesting, because the thrust-to-weight ratio of its default quadrotor is $\approx 12$ (cf. Figure 4) and hence $2\times$ the upper limit of $5$ of our domain randomization range. We find that the z-error is significant, but the foundation policy still controls it robustly and can track agile trajectories. This shows the remarkable robustness and out-of-distribution generalization of the foundation policy resulting from the RAPTOR method.
3 Discussion
In our extensive experimental evaluations, we find that our proposed RAPTOR framework produces a highly robust and versatile foundation policy that can control a broad range of quadrotors in a large variety of situations. We believe that our experiments validate the design choices in the RAPTOR architecture.
To facilitate research, we aim for the highest levels of reproducibility, publishing all training code for the benefit of the community. Furthermore, we make it extremely easy to integrate the foundation policy into existing flight controller software and simulators, such that the barrier to use it as a baseline is minimal.
4 Materials and Methods
Figure 7: (A) A probabilistic graphical model (Bayesian Network) of the dynamics and control of a random quadrotor. This formal model allows us to derive the RAPTOR architecture from probabilistic principles. (B): Foundation policy network architecture. (C): Illustration of inferring dynamics parameters by reasoning about the observed input/output behavior of the system.
@@ 4 Materials and Methods
We formulate the quadrotor control problem as a Bayes Adaptive Partially Observable Markov Decision Process (BAPOMDP)  (?) defined by the tuple $(\mathcal{S},\mathcal{S}_{0},\mathcal{D},\mathcal{A},\mathcal{T},r,\boldsymbol{\Omega},\mathbf{o},\gamma,\boldsymbol{\Xi})$. $\mathcal{S}$ is the set of states $\mathbf{s}=\{\mathbf{p},\mathbf{q},\mathbf{v},\boldsymbol{\omega},\mathbf{a}_{t-1},\boldsymbol{\omega}_{m}\}$ consisting of position, orientation, linear/angular velocity, previous action and motor states respectively. The previous action is included in the state because the reward function penalizes the change in action. $\mathcal{S}_{0}$ is the initial state distribution with position, orientation, linear/angular velocity uniformly sampled up to $10\cdot l_{\text{arm}}$, $90^{\circ}$, $1$ m/s, $1$ rad/s respectively. With a probability of $10\%$, the initial state is overwritten with the target state (all zeros). $\mathcal{D}$ is the termination relation that includes all states outside $20\cdot l_{\text{arm}}$ m, $2$ m/s, $35$ rad/s (position, linear/angular velocity respectively). $\mathcal{A}$ is the set of actions $\mathbf{a}=\{\omega_{\text{sp}_{0}},\omega_{\text{sp}_{1}},\omega_{\text{sp}_{2}},\omega_{\text{sp}_{3}}\}$ (individual motor commands). The transition probabilities $\mathcal{T}$ are defined as $\mathrm{p}(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t},\boldsymbol{\Xi})$ and implemented by the L2F simulator  (?). The reward function is deterministic:
$\mathrm{r}(\mathbf{s}_{t},\mathbf{a},\mathbf{s}_{t+1})=\|\mathbf{p}\|_{2}+0.2\cdot\arccos(1-|q_{z}|)+\|\mathbf{a}_{t}-\mathbf{a}_{t-1}\|_{2}+1.5-100\cdot\mathbf{1}[\mathrm{terminal}(\mathbf{s}_{t+1})]$ (1)
@@
We include the next state in the reward function to be able to inflict the termination penalty. The observation space $\mathcal{O}$ contains a subset of the state space $\mathbf{o}=\{\mathbf{p},\mathbf{R}(\mathbf{q}),\mathbf{v},\boldsymbol{\omega},\mathbf{a}_{t-1}\}$, occluding the motor states $\boldsymbol{\omega_{m}}$ (not observable on most real-world platforms). $\gamma=0.99$ is the discount factor and the domain parameters are collected in $\boldsymbol{\Xi}=\{m,l_{\text{arm}},c_{f_{0}},c_{f_{1}},c_{f_{2}},c_{m},\mathbf{J}=\text{diag}(J_{xx},J_{yy},J_{zz}),T_{m\uparrow},T_{m\downarrow}\}$ containing the mass, arm length, thrust-curve coefficients (zeroth, first and second order), moment coefficient, inertia matrix and rising/falling edge motor delays respectively. The difference between a BAPOMDP and a POMDP is that we can factor out the system parameters $\boldsymbol{\Xi}$ (which would have to be encoded into the state in a normal POMDP). This factorization allows us to implement the inductive bias that the parameters are only sampled once at the beginning of the episode and remain constant throughout.
@@
Using this BAPOMDP framework, in the following, we provide a formal derivation of our method using a probabilistic graphical model  (?). The full model is shown in Figure 7A. The goal is to model the decision-making at time $t$. At timestep $t$, $t$ previous (and current) observations have been observed (shaded nodes) as well as t1t-1 previous actions. We explicitly place minimal assumptions on the policy that decides the previous actions, which shows by the previous actions being causally fully connected to all previous actions and observations. By causally fully connected, we mean that an action at timestep $t$ might depend on observations $\mathbf{o}_{0},\cdots,\mathbf{o}_{t}$ and actions $\mathbf{a}_{0},\cdots,\mathbf{a}_{t-1}$. Hence, the only assumption is that the policy generating the previous trajectory is causal.
@@
Using d-Separation  (?), an algorithm to prove conditional independencies in probabilistic graphical models, we can show that not observing $\mathbf{s}_{t}$ and $\boldsymbol{\Xi}$ makes the distribution over the optimal action $\mathbf{a}_{t}^{*}$ dependent on other variables that could carry information about them. The most direct example is $\mathbf{o}_{t}$. Since $\mathbf{o}_{t}$ is derived from $\mathbf{s}_{t}$, it can carry information about $\mathbf{s}_{t}$ and in our case it carries almost all the information about the state, apart from the motor speeds. Hence, when $\mathbf{o}_{t}$ is observed, this changes the distribution over the optimal action $\mathbf{a}_{t}^{*}$. The same argument holds for $\mathbf{o}_{t-1}$ and earlier observations because, e.g., $\mathbf{o}_{t-1}$ carries information about $\mathbf{s}_{t-1}$ and $\mathbf{s}_{t-1}$ (in combination with the observable $\mathbf{a}_{t-1}$) carries all the information to infer the distribution over $\mathbf{s}_{t}$. This is because $\mathbf{s}_{t-1}$, $\mathbf{a}_{t-1}$, $r_{t-1}$, and $\boldsymbol{\Xi}$ form the causal Markov blanket  (?) for $\mathbf{s}_{t}$. For the causal Markov blanket, random variables that lie in the future are removed. Therefore, using the d-Separation rules, we can see that the distribution over the optimal action $\mathbf{a}_{t}^{*}$ is dependent on all previous observations and actions. There is no set of observable variables that could form a Markov blanket and shield this dependency. This motivates our design decision that the foundation policy, which is the student policy from the Meta-Imitation Learning perspective, is dependent on all previous observations and actions. Intuitively, the dependence on all previous observations and actions makes sense because any of the individual observations, or, more likely, the combination of observations over time, contains information about the dynamics parameters $\boldsymbol{\Xi}$ and about the ground-truth state $\mathbf{s}$.
This is illustrated in Figure 3 and Figure 7C where we show how the relation of observations over time carries information about the dynamics parameters. By looking at the observation-action history, the policy can observe that e.g., the effect of the RPM setpoint actions is always delayed by one timestep. Note that this is a simplification by using a pure delay/dead time, while in reality the temporal relationship between RPM setpoints and motor speeds is more closely modeled as a first-order low-pass. Furthermore, it can be observed that an RPM setpoint of $1$ corresponds to an observed acceleration of $1$, but an RPM setpoint of $2$ corresponds to an observed acceleration of $4$. Hence it can infer the curvature of the thrust curve.
@@
Instead, we want to model it as closely as possible and hence phrase the problem as variational inference, where we try to minimize the Kullback-Leibler (KL) divergence between the predicted optimal action distribution and the actual optimal action distribution. Note that due to the non-linearities in the system dynamics, etc., the actual optimal action distribution is not tractable and that we approximate it by training expert teacher policies until convergence.
@@
While the Bayesian Network in Figure 7A is a formal/mathematical probabilistic model, Figure 1B shows our proposed practical method for modeling the various conditional probability distributions that constitute it:
@@ 4.1 Domain Randomization
Please refer to the Supplementary Material  (?) for the detailed equations that establish the physically plausible correlations.
4.2 Training Methodology
@@
4.2.1 Pre-Training
In the pre-training phase, we take $1000$ quadrotors sampled from the distribution described in Section 4.1 and train a dedicated expert policy for each of them by creating an independent MDP for each set of sampled domain parameters $\boldsymbol{\Xi}$. We also make the ground-truth states directly observable because this expert policy does not need to be deployed on hardware. Since we are not constrained by the deployment onto hardware, we can overparameterize it to aid the training  (?). We use three layers with a hidden dimensionality of $64$, making each teacher policy $3\times$ larger than the condensed foundation policy in terms of parameters. The training pipeline is adapted from  (?) with five modifications:
@@ 4.2.1 Pre-Training
    Switching from TD3 to SAC because we observed slightly more robust training dynamics in SAC.
@@
We do these modifications to trade off wall-clock training time for highly reliable training dynamics, and we found them to yield high-quality teacher policies in all $1000$ cases without cherry-picking random seeds and without requiring case-by-case modifications, even though the dynamics of the quadrotors vary drastically as described in Section 4.1.
The teachers observe the states fully $\mathbf{o}_{\text{teacher}}=\{\mathbf{p},\mathbf{R}(\mathbf{q}),\mathbf{v},\boldsymbol{\omega},\mathbf{a}_{t-1},\boldsymbol{\omega}_{m}\}$.
We train using a position control objective where the quadrotor is initialized in a random state (e.g., up to $90^{\circ}$ tilt) and the policy has to navigate it back to the origin with zero yaw while also minimizing linear/angular velocity and action changes. To prepare the policy for trajectory tracking, we also train with the objective of tracking a relatively slow trajectory sampled from a second-order Langevin process. The Supplementary Materials  (?) contain additonal details about the motivation for and the implementation of the distribution over reference trajectories.
4.2.2 Meta-Imitation Learning
@@
After training $1000$ teacher policies, we would like to distill all of their behaviors into a single student foundation policy. From the perspective of each teacher policy, there are no hidden/latent parameters because each teacher policy can assume that it always interacts with the same quadrotor $\boldsymbol{\Xi}$. However, in aggregate, and from the perspective of the student policy, the parameters of the quadrotor $\boldsymbol{\Xi}$ it is interacting with are not observable and need to be inferred as described in Section 4. Due to the variable number of past steps (cf. Figure 7A), we design a Gated Recurrent Unit (GRU)  (?)-based foundation policy architecture as displayed in Figure 7B.
The relatively small hidden dimensionality of $16$ is justified by the scaling experiments in Section 2.3. Due to the recurrence, the policy can theoretically ”access” all the previous observations and actions.
We refer to our proposed algorithm as Meta-Imitation Learning because the student not only needs to learn to recreate the teachers’ outputs from a different set of inputs/sensors, but also must learn to perform inference about the current MDP that it is acting in. Each quadrotor constitutes a separate MDP because the transition function varies based on the random dynamics parameters $\boldsymbol{\Xi}$. We currently only tackle the case where the dynamics parameters change, but in the future, we will also incorporate, e.g., changes in the reward function.
Our proposed method is conceptually similar to the DAgger algorithm  (?) but differs in two key ways, firstly in the aforementioned requirement for meta-learning and secondly in that we perform on-policy data collection (after warm-up) and learning while DAgger is strictly off-policy. Figure 8 shows our full proposed algorithm consisting of the sampling of $1000$ random quadrotors and the two main learning phases: pre- and post-training.
@@
4.3 Computational Aspects
Computationally, the separation into pre-training and post-training/meta-imitation learning is a major advantage of the RAPTOR framework. This decouples the time/compute-intensive pre-training and renders it ”embarrassingly parallel”  (?). This can be seen from Figure 1B, where the teacher training processes are independent. Hence, we can horizontally scale out the number of training processes over multiple processors and/or machines, linearly speeding the pre-training up to the ceiling, which is the $31$ m duration of each training run (at $1000$ cores in parallel). In contrast to LLM pre-training, where the gradients have to be communicated between all nodes at every training step, in RAPTOR, pre-training is entirely independent and communication/synchronization is only required after pre-training, in the distillation/meta-imitation learning phase, which is about three orders of magnitude less computationally intensive than the pre-training.

### v2

@@ Results
Disturbances
Figure 6: RAPTOR Policy in Different Situations. (A): The foundation policy is activated in mid-air, starting with a linear velocity of $4.5$ m/s. (B): Crazyflie Brushless tracking an agile trajectory. (C): Long exposure photo of three consecutive loops of the Crazyflie Brushless trajectory. (D): The foundation policy recovers from being poked and tilting $90^{\circ}$. (E): A tool is rested onto the quadrotor flown by the foundation policy. (F): A quadrotor with four different propellers (2- and 3-blade) is tracking a trajectory using the foundation policy.
@@
Additionally, we also test swapping out $1$, 22, and $3$ propellers with random three-blade ones (the original ones are two-blade) and find that, in any of these configurations, the foundation policy still stably controls the vehicle and is even able to track trajectories (see Figure 6F and Movie S6). Note that the distribution of quadrotors described in the Domain Randomization subsection covers only quadrotors with identical thrust curves for all four motors/propellers. Hence, neither the teacher policies nor the foundation policy has ever experienced differing thrust curves on the same quadrotor. Yet, we find that the foundation policy can generalize zero-shot outside of this distribution, and control the quadrotor well.
In the absence of disturbances, we find that the foundation policy yields a remarkably repeatable performance, as can be seen from the trajectories in Figure 5, which show $5$ consecutive full loops each. This also answers the posed research question about the trajectory length generalization of the foundation policy. During Meta-Imitation Learning, we train the policy with sequences of $500$ steps, corresponding to $5$ s of flight time. During inference, we test the foundation policy with $5$ consecutive iterations of, for example, a $10$ s, amounting to a trajectory length of $50$ s. This shows a $10\times$ context window size extrapolation.
@@
To further test the out-of-distribution generalization beyond the aforementioned real-world tests including flexible frame, mixed propellers, and context window extrapolation, we test simulation-to-simulation transfer to the Flightmare simulator. This transfer is interesting, because the thrust-to-weight ratio of its default quadrotor is $\approx 12$ (see Figure 4) and hence $2\times$ the upper limit of $5$ of our domain randomization range. We find that the z-error is substantial, but the foundation policy still controls it robustly and can track agile trajectories. This shows the remarkable robustness and out-of-distribution generalization of the foundation policy resulting from the RAPTOR method.
Discussion
In our extensive experimental evaluations, we find that our proposed RAPTOR framework produces a highly robust and versatile foundation policy that can control a broad range of quadrotors in a large variety of situations. Compared to state-of-the-art solutions for end-to-end quadrotor control, our method is able to adapt to a wide range of physical platforms without re-training, through in-context learning. Even though existing methods rely on overfitting a control policy for each platform, our method is able to approximately match their performance while providing the flexibility of zero-shot adaptation and being substantially more light-weight, computationally. This allows, for example, deployment on platforms whose parameters have not yet been identified. Due to the minimal compute requirements of the tiny resulting policy, it can be deployed even on the smallest quadrotors. Hence, we believe that our experiments validate the design choices in the RAPTOR architecture.
Nevertheless, through experimentation in the real world, we find that there are many avenues for future research based on our proposed architecture, including the following $4$ directions:
  1. 1.
    Simulation-to-reality gap in flight controller firmware implementations: We find that a large part of the simulation-to-reality gap comes from delays in the flight controller firmware and infrastructure around the foundation policy as well as the state estimation. We see potential for introducing more domain randomization to simulate delays that appear during deployment, as well as moving to a fully end-to-end architecture by cutting out the state estimation and directly feeding IMU data into the policy.
  2. 2.
    Over-reliance on linear velocity observations: We find that the foundation policy overindexes on the linear velocity observations to perform emergent system identification. In real-world deployments, there is usually a delay from the infrastructure (motion capture system) or other disturbing factors (such as drift in GPS). In future work, we believe this problem can be alleviated by adding more observation noise to the linear velocity observations and adding accelerometer measurements to the observation space.
  3. 3.
    Limited domain randomization range: Although our distribution over quadrotor dynamics already covers a vast range of quadrotors (and arguably $50\%$ of quadrotors that are being used in the real world), our experiments showed that there are limits to the out-of-distribution generalization. We believe that the lower tracking performance on the Flightmare platform (thrust-to-weight ratio of $\approx 12$ vs. $\leq 5$ during training) can be alleviated by extending the domain randomization range because when limiting its thrust-to-weight ratio to $5$, we find it to yield excellent tracking performance.
  4. 4.
    Trajectory tracking lookahead: We find that lookahead-free trajectory tracking is posing a limit and believe that integrating lookahead into the RAPTOR architecture can substantially improve the trajectory tracking performance.
Materials and Methods
@@ Discussion
We formulate the quadrotor control problem as a Bayes Adaptive Partially Observable Markov Decision Process (BAPOMDP) (?) defined by the tuple $(\mathcal{S},\mathcal{S}_{0},\mathcal{D},\mathcal{A},\mathcal{T},\mathrm{r},\mathcal{O},\mathbf{o},\gamma,\boldsymbol{\Xi})$. $\mathcal{S}$ is the set of states $\mathbf{s}=\{\mathbf{p},\mathbf{q},\mathbf{v},\boldsymbol{\omega},\mathbf{a}_{t-1},\boldsymbol{\omega}_{m},\mathbf{f}_{\text{ext}}\}$ consisting of position, orientation, linear/angular velocity, previous action, motor states and a random external force, respectively. The previous action is included in the state because the reward function penalizes the change in action. $\mathcal{S}_{0}$ is the initial state distribution with position, orientation, linear/angular velocity uniformly sampled up to $10\cdot l_{\text{arm}}$, $90^{\circ}$, $1$ m/s, $1$ rad/s, respectively. With a probability of $10\%$, the initial state is overwritten with the target state (all zeros). The random force $\mathbf{f}_{\text{ext}}$ resembles, for example, wind disturbances and is sampled from a zero-mean normal distribution with a standard deviation derived from the sampled dynamics parameters (see supplementary materials (?)). $\mathcal{D}$ is the termination relation that includes all states outside $20\cdot l_{\text{arm}}$ m, $2$ m/s, $35$ rad/s (position, linear/angular velocity respectively). $\mathcal{A}$ is the set of actions $\mathbf{a}=\{\omega_{\text{sp}_{0}},\omega_{\text{sp}_{1}},\omega_{\text{sp}_{2}},\omega_{\text{sp}_{3}}\}$ (individual motor commands). The transition probabilities $\mathcal{T}$ are defined as $\mathrm{p}(\mathbf{s}_{t+1}|\mathbf{s}_{t},\mathbf{a}_{t},\boldsymbol{\Xi})$ and implemented by the L2F simulator (?). The reward function is deterministic:
$\mathrm{r}(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1})=-\|\mathbf{p}\|_{2}-0.2\cdot\arccos(1-|q_{z}|)-\|\mathbf{a}_{t}-\mathbf{a}_{t-1}\|_{2}+1.5-100\cdot\mathbf{1}[\mathrm{terminal}(\mathbf{s}_{t+1})]$ (1)
@@
We include the next state in the reward function to be able to inflict the termination penalty. The observation space $\mathcal{O}$ contains a subset of the state space $\mathbf{o}=\{\mathbf{p},\mathbf{R}(\mathbf{q}),\mathbf{v},\boldsymbol{\omega},\mathbf{a}_{t-1}\}$, occluding the motor states $\boldsymbol{\omega_{m}}$ (not observable on most real-world platforms) and the external disturbance. $\gamma=0.99$ is the discount factor and the domain parameters are collected in $\boldsymbol{\Xi}=\{m,l_{\text{arm}},c_{f_{0}},c_{f_{1}},c_{f_{2}},c_{m},\mathbf{J}=\text{diag}(J_{xx},J_{yy},J_{zz}),T_{m\uparrow},T_{m\downarrow}\}$ containing the mass, arm length, thrust-curve coefficients (zeroth, first and second order), moment coefficient, inertia matrix and rising/falling edge motor delays respectively. The difference between a BAPOMDP and a POMDP is that we can factor out the system parameters $\boldsymbol{\Xi}$ (which would have to be encoded into the state in a normal POMDP). This factorization allows us to implement the inductive bias that the parameters are only sampled once at the beginning of the episode and remain constant throughout.
@@
Using this BAPOMDP framework, in the following, we provide a formal derivation of our method using a probabilistic graphical model (?). The full model is shown in Figure 7A. The goal is to model the decision-making at time $t$. At timestep $t$, $t$ previous observations and actions, as well as the current observation have been observed (shaded nodes). We explicitly place minimal assumptions on the policy that decides the previous actions, which shows by the previous actions being causally fully connected to all previous actions and observations. By causally fully connected, we mean that an action at timestep $t$ might depend on observations $\mathbf{o}_{0},\cdots,\mathbf{o}_{t}$ and actions $\mathbf{a}_{0},\cdots,\mathbf{a}_{t-1}$. Hence, the only assumption is that the policy generating the previous trajectory is causal.
Figure 7: (A): A probabilistic graphical model (Bayesian Network) of the dynamics and control of a random quadrotor. This formal model allows us to derive the RAPTOR architecture from probabilistic principles. (B): Foundation policy network architecture. (C): Illustration of inferring dynamics parameters by reasoning about the observed input/output behavior of the system.
@@
Using d-Separation (?), an algorithm to prove conditional independencies in probabilistic graphical models, we can show that not observing $\mathbf{s}_{t}$ and $\boldsymbol{\Xi}$ makes the distribution over the optimal action $\mathbf{a}_{t}^{*}$ dependent on other variables that could carry information about them. The most direct example is $\mathbf{o}_{t}$. Since $\mathbf{o}_{t}$ is derived from $\mathbf{s}_{t}$, it can carry information about $\mathbf{s}_{t}$ and in our case it carries almost all the information about the state, apart from the motor speeds. Hence, when $\mathbf{o}_{t}$ is observed, this changes the distribution over the optimal action $\mathbf{a}_{t}^{*}$. The same argument holds for $\mathbf{o}_{t-1}$ and earlier observations because, for example, $\mathbf{o}_{t-1}$ carries information about $\mathbf{s}_{t-1}$ and $\mathbf{s}_{t-1}$ (in combination with the observable $\mathbf{a}_{t-1}$) carries all the information to infer the distribution over $\mathbf{s}_{t}$. This is because $\mathbf{s}_{t-1}$, $\mathbf{a}_{t-1}$, $r_{t-1}$, and $\boldsymbol{\Xi}$ form the causal Markov blanket (?) for $\mathbf{s}_{t}$. For the causal Markov blanket, random variables that lie in the future are removed. Therefore, using the d-Separation rules, we can see that the distribution over the optimal action $\mathbf{a}_{t}^{*}$ is dependent on all previous observations and actions. There is no set of observable variables that could form a Markov blanket and shield this dependency. This motivates our design decision that the foundation policy, which is the student policy from the Meta-Imitation Learning perspective, is dependent on all previous observations and actions. Intuitively, the dependence on all previous observations and actions makes sense because any of the individual observations, or, more likely, the combination of observations over time, contains information about the dynamics parameters $\boldsymbol{\Xi}$ and about the ground-truth state $\mathbf{s}$.
This is illustrated in Figure 3 and Figure 7C where we show how the relation of observations over time carries information about the dynamics parameters. By looking at the observation-action history, the policy can observe that, for example, the effect of the motor commands is always delayed by one timestep. Note that this is a simplification by using a pure delay/dead time, whereas in reality the temporal relationship between motor commands and motor speeds is more closely modeled as a first-order low-pass filter. Furthermore, it can be observed that a motor command of $1$ corresponds to an observed acceleration of $1$, but a motor command of $2$ corresponds to an observed acceleration of $4$. Hence it can infer the curvature of the thrust curve.
@@
Instead, we want to model it as closely as possible and hence phrase the problem as variational inference, where we try to minimize the Kullback-Leibler (KL) divergence between the predicted optimal action distribution and the actual optimal action distribution. Note that due to the non-linearities in the system dynamics, the actual optimal action distribution is not tractable and that we approximate it by training expert teacher policies until convergence.
@@
Whereas the Bayesian Network in Figure 7A is a formal/mathematical probabilistic model, Figure 1B shows our proposed practical method for modeling the various conditional probability distributions that constitute it:
@@
Please refer to the Supplementary Materials (?) for the detailed equations that establish the physically plausible correlations.
Training Methodology
@@
Pre-Training
In the pre-training phase, we take $1000$ quadrotors sampled from the distribution described in the Domain Randomization subsection and train a dedicated expert policy for each of them by creating an independent MDP for each set of sampled domain parameters $\boldsymbol{\Xi}$. We also make the ground-truth states directly observable because this expert policy does not need to be deployed on hardware. Since we are not constrained by the deployment onto hardware, we can overparameterize it to aid the training (?). We use fully-connected three-layer neural networks with a hidden dimensionality of $64$, making each teacher policy $3\times$ larger than the condensed foundation policy in terms of parameters. The training pipeline is adapted from (?) with five modifications:
@@
    Switching from TD3 to SAC because we observed slightly more robust training dynamics in SAC. Note that we use an off-policy method because in our experience they are much more reliable than on-policy methods like PPO and, for example, do not require cherry-picking random seeds. The latter is of utmost importance because we do not have the capacity to supervise each of the $1000$ training runs, and hence rely on each of them finding a good policy without any adjustments to hyperparameters.
@@
We do these modifications to trade off wall-clock training time for highly reliable training dynamics, and we found them to yield high-quality teacher policies in all $1000$ cases without cherry-picking random seeds and without requiring case-by-case modifications, even though the dynamics of the quadrotors vary drastically as described in the Domain Randomization subsection.
The teachers observe the states fully $\mathbf{o}_{\text{teacher}}=\{\mathbf{p},\mathbf{R}(\mathbf{q}),\mathbf{v},\boldsymbol{\omega},\mathbf{a}_{t-1},\boldsymbol{\omega}_{m},\mathbf{f}_{\text{ext}}\}$.
We train using a position control objective where the quadrotor is initialized in a random state (for example, up to $90^{\circ}$ tilt) and the policy has to navigate it back to the origin with zero yaw while also minimizing linear/angular velocity and action changes. To prepare the policy for trajectory tracking, we also train with the objective of tracking a relatively slow trajectory sampled from a second-order Langevin process. The Supplementary Materials (?) contain additional details about the motivation for and the implementation of the distribution over reference trajectories.
Meta-Imitation Learning
After training $1000$ teacher policies, we would like to distill all of their behaviors into a single student foundation policy. From the perspective of each teacher policy, there are no hidden/latent parameters because each teacher policy can assume that it always interacts with the same quadrotor $\boldsymbol{\Xi}$. However, in aggregate, and from the perspective of the student policy, the parameters of the quadrotor $\boldsymbol{\Xi}$ it is interacting with are not observable and need to be inferred as described in the Materials and Methods section. Due to the variable number of past steps (see Figure 7A), we design a Gated Recurrent Unit (GRU) (?)-based foundation policy architecture as displayed in Figure 7B.
The relatively small hidden dimensionality of $16$ is justified by the scaling experiments in the Meta-Imitation Learning subsection of the Results section. Due to the recurrence, the policy can theoretically ”access” all the previous observations and actions.
We refer to our proposed algorithm as Meta-Imitation Learning because the student not only needs to learn to recreate the teachers’ outputs from a different set of inputs/sensors, but also must learn to perform inference about the current MDP that it is acting in. Each quadrotor constitutes a separate MDP because the transition function varies based on the random dynamics parameters $\boldsymbol{\Xi}$. We currently only tackle the case where the dynamics parameters change, but in the future, we will also incorporate, for example, changes in the reward function.
Our proposed method is conceptually similar to the DAgger algorithm (?) but differs in two key ways, firstly in the aforementioned requirement for meta-learning and secondly in that we perform on-policy data collection (after warm-up) and learning whereas DAgger is performing on-policy data collection but also uses off-policy data (the aggregated dataset) for learning. Figure 8 shows our full proposed algorithm consisting of the sampling of $1000$ random quadrotors and the two main learning phases: pre- and post-training.
@@
Computational Aspects
Computationally, the separation into pre-training and post-training/meta-imitation learning is a major advantage of the RAPTOR framework. This decouples the time/compute-intensive pre-training and renders it ”embarrassingly parallel” (?). This can be seen from Figure 1B, where the teacher training processes are independent. Hence, we can horizontally scale out the number of training processes over multiple processors and/or machines, linearly speeding the pre-training up to the ceiling, which is the $31$ min duration of each training run (at $1000$ cores in parallel). In contrast to LLM pre-training, where the gradients have to be communicated between all nodes at every training step, in RAPTOR, pre-training is entirely independent and communication/synchronization is only required after pre-training, in the distillation/meta-imitation learning phase, which is about three orders of magnitude less computationally intensive than the pre-training.